OpenStudy (anonymous):

If f(x) = x + 3 and g(x) = −4x − 1, find g[f(4)].

OpenStudy (anonymous):

\(f(4)=4+3=7\) \(g(7)=?\)

OpenStudy (anonymous):

g[f(4)]=-4*7-1=-29